OpenStudy (anonymous):

Find the vertex, focus, directrix, and Latus Rectum endpoint. x^2=8y Would you set the Y by itself first and then proceed with finding the rest?

OpenStudy (anonymous):

i found it its u cutie lol jk

OpenStudy (anonymous):

Thank you, but would I begin by setting the Y by itself and find the rest from there?

OpenStudy (aum):

x^2=8y y = 1/8 * x^2 Compare this to the vertex form of the parabola: y = a * (x-h)^2 + k where (h,k) is the vertex. What is the vertex of the given parabola?

OpenStudy (anonymous):

Hmm...(-1,-1/8)

OpenStudy (aum):

Compare y = 1/8 * x^2 and y = a * (x-h)^2 + k What is a, h and k?

OpenStudy (anonymous):

x^2 is h and 1/8 is K?

OpenStudy (aum):

Compare y = 1/8 * x^2 and y = a * (x-h)^2 + k a = 1/8 h = 0 k = 0 Vertex is (h,k) and for this parabola vertex is (0,0). The distance 'p' between the vertex and the focus which is also the distance between the vertex and the directrix is given by p = 1/(4a). If a = 1/8 What is p?

OpenStudy (aum):